OpenStudy (anonymous):

Can figuring out the slope of a line get any simpler? I'm having a really hard time even getting started. I'm so not a math person.

OpenStudy (anonymous):

slope is change in y over change in x

OpenStudy (anonymous):

Basically I pick two points and give the difference but in a formula?

OpenStudy (anonymous):

Do you know a REALLY GOOD website where I can get an example? I have an assignment due tomorrow and could use some help. Thanks!

OpenStudy (anonymous):

lol this website can't support multiple tabs. ignore that answer :P

OpenStudy (anonymous):

programming of this website is bad.. but if you have two points , lets call them A and B. A (6,8) and B (5,4). the slope is simply (y2 - y1) / (x2 - x1) = (4-8)/(5-6) = 4 The answer is the same if you invert the order of 1 and 2.